OpenStudy (anonymous):

integral of (x^4)/(square root of (2x^5-x^10))

OpenStudy (lgbasallote):

\[\large \int \frac{x^4}{\sqrt{x^5(2 - x^2)}}\] does that help?

OpenStudy (anonymous):

yeah i got that far but cant figure it out

OpenStudy (lgbasallote):

\[\large \int \frac{x^4}{x^4 \sqrt{x(2-x^2)}}\]

OpenStudy (anonymous):

thats not right

OpenStudy (lgbasallote):

oops that's x^2 in the denominator lol \[\large \int \frac{x^4}{x^2 \sqrt{x(2 - x^2)}}\] sorr

OpenStudy (anonymous):

how is that the same as the original question?

sam (.sam.):

\[\Huge \int\limits_{}^{}\frac{x^4}{\sqrt{2x^5-x^{10}}}\]

OpenStudy (mani_jha):

Taking \[x ^{5}=z\] \[5x ^{4}dx=dz\] The expression becomes: \[\int\limits_{}^{} dz/(5\sqrt{2z-z ^{2}})\] Can you proceed from here?

OpenStudy (anonymous):

amazing @Mani_Jha

sam (.sam.):

Somehow I find wolfram can't do simple maths http://www.wolframalpha.com/input/?i=integrate+%28x%5E4%29%2Fsqrt%282x%5E5-x%5E%2810%29%29 :/

OpenStudy (anonymous):

This integral does not have closed form answer. Are you sure you typed it right?